OpenStudy (anonymous):

Factor the following expression: 6(x^2 - 4x + 4)^2 + (x^2 - 4x + 4) -1 So I simplfied it to: 6x^2 + x -1 I use the crisscross method which is REALLY fast and accurate.... but I couldn't find an answer... Can you tell me a hint? :D

OpenStudy (anonymous):

Nevermind got it.

OpenStudy (anonymous):

Thanks anyways :D

OpenStudy (campbell_st):

this is a simple way let \[u = x^2 -4x + 4\] using this substitution you would get a quadratic \[6u^2 + u - 1\] which can be factored once its factored then you can replace u by the original substitution

OpenStudy (anonymous):

I get that part i was stuck on finding the numbers that x to -6 and add to +1

OpenStudy (campbell_st):

ok... if you have \[ax^2 + bx + c\] I always do \[a \times c\] and find the factors that add to b so 6 x - 1 = -6 what factors of -6 add to 1

OpenStudy (campbell_st):

try 3 ans - 2

OpenStudy (anonymous):

so \[[3(x^2-4x+4) -2][2(x^2-4x+4)+3]\]?

OpenStudy (anonymous):

Answers say I'm wrong... it says \[(2x^2 - 8x + 9)(3x^2 - 12x + 11)\]

OpenStudy (campbell_st):

not quite right... this is the method I use when I have the factors correct...  next remove the common factors from each set of brackets